Which of the following pairs represent the same rational number?
(-2/-3) and (2/3)

Grade:12th pass

1 Answers

Harshit Singh
2 years ago
Dear Student

We have to check the given pair represents the same rational number.
Then,
-2/-3 = 2/3
2/3= 2/3∵2/3 = 2/3
∴-2/-3 = 2/3
So, the given pair is represents the same rational number.

Thanks